# New black hole mergers in the LIGO–Virgo O3 data from a gravitational wave search including higher-order harmonics

### v1

@@ Abstract
Nearly all of the previous gravitational wave (GW) searches in the LIGO–Virgo data included GW waveforms with only the dominant quadrupole harmonic, i.e., omitting higher-order harmonics which are predicted by general relativity. Based on the techniques developed in Wadekar et al. (2023a, b), we improve the IAS pipeline by () introducing higher harmonics in the GW templates, () downweighting noise transients (“glitches”) to improve the search sensitivity to high-mass and high-redshift binary black hole (BBH) mergers. We find 14 new candidate BBH mergers with on running our pipeline over the public LIGO–Virgo data from the O3 run (we use the detection threshold as the astrophysical probability, , being over 0.5, following the approach of other pipelines). We also broadly recover the high-significance events from earlier catalogs, except some which were either vetoed or fell below our SNR threshold for trigger collection.
@@ I Introduction
Gravitational waves (GW) from roughly 100 binary black hole (BBH) mergers have been detected in the data from the first three observing runs of the LIGO–Virgo–Kagra collaboration (LVK), combining detections by the LVK Abbott et al. (2016a, 2019, 2021); The LIGO Scientific Collaboration et al. (2021); Abbott et al. (2021) and by other independent pipelines Venumadhav et al. (2019, 2020); Olsen et al. (2022); Nitz et al. (2019, 2020, 2021, 2021); Chia et al. (2023) who access the data made publicly available through the Gravitational Wave Open Science Center (GWOSC) Vallisneri et al. (2015). Nearly all the previous template-bank based searches have approximated GW waveforms to have only the dominant quadrupole mode, , where the frequency of GWs is twice the orbital frequency.
@@
The mode appears at the lowest order in post-Newtonian (PN) expansion, but there are a number of higher harmonics originating from the subsequent orders of the PN expansion (e.g., the mode has thrice the orbital frequency) Thorne (1980); Pan et al. (2011). Approximate formulae for the relative strength of frequency-domain amplitudes of two of the higher-order modes (HM) are given by Wadekar et al. (2023a)
@@
where is the mass ratio (), is the orbital frequency and is the binary inclination ( corresponds to the edge-on configuration). These harmonics have a number of important applications for characterizing properties of BBHs. If one uses the mode waveform alone in inferring the BBH properties, we have the following important degeneracies between different binary parameters: luminosity distance and ; and effective spin (); polarization and initial orbital phase (). All three degeneracies are broken by HM Mills and Fairhurst (2021); Varma et al. (2014). This follows from the amplitudes of HM having a different dependence on and as compared to that of the mode (see Eq. (1)), and similarly the phases of HM have different dependence on . It is worth mentioning that the degeneracy between distance and inclination is currently one of the most important limiting factors in inferring the Hubble parameter from GW systems and HM can be crucial in breaking this degeneracy. Similarly, events which have HM also have the best measured mass ratios (e.g., Abbott et al. (2020a, b)). HM become important in the strong gravity regime close to merger and are thus useful to test the multipolar structure of gravity Kastha et al. (2018, 2019); Dhanpal et al. (2019); Islam et al. (2020); Capano et al. (2021), and to study properties of the merger remnant such as the recoil kick.
@@
Multiple waveform models including HM have been recently made available London et al. (2018); García-Quirós et al. (2020); Mehta et al. (2017), and they are currently being used ubiquitously in parameter estimation studies. However, HM are still not included in current search pipelines (as of now, the only exception is Ref. Chandra et al. (2022), which searched for binaries with nearly edge-on inclinations). The reason behind this is that for making template banks with the mode, one only needs to sample over the masses and spins of black holes (). However, for constructing banks including HM, one needs to additionally sample over the inclination and initial phase (, ). These additional degrees of freedom can cause: the size of the template banks and the cost of search to increase significantly (by a factor of 100 for the case of HM Harry et al. (2018); Chandra et al. (2022); Schmidt et al. (2023)), loss of sensitivity in the search as there is increased number of background triggers Capano et al. (2014). To tackle the first issue, we found more efficient ways of including HM in our template banks (which we presented in detail in our first companion paper Wadekar et al. (2023a)). Regarding the second issue, we devise new detection statistics which includes marginalization over the amplitudes and phases of higher modes in our second companion paper Wadekar et al. (2023b). We present an overview of these new techniques in Section II.
@@
We see from Eq. (1) that, at a given orbital frequency, the relative HM amplitudes increase as a power-law with . Moreover, due to the shape of the current detector power spectral density (PSD), the signal-to-noise ratio (SNR) contribution of HM is additionally boosted at high masses and redshifts. In this regime, the mode can cut off below the detector’s sensitive band (the detector PSD sharply increases below Hz) and HM, being at higher frequencies, naturally have an increased contribution to the SNR. The background corresponding to the high-mass regime is however heavily dominated by short-duration noise transients (called instrumental “glitches”). This is because the high-mass waveforms have a very short duration in band, which makes it easy for transients like blip glitches to mimic real signals. We therefore introduce additional methods in the IAS pipeline to downweight the background at high masses. These methods are detailed in our second companion paper Wadekar et al. (2023b) and we present an overview in Section II.3.
@@ II Pipeline improvements
Here, we present an outline of the additions that we made to the -only IAS pipeline Olsen et al. (2022); Venumadhav et al. (2019).
@@ II.1 Adding HM in template banks
In this sub-section, we summarize the methodology for constructing template banks for HM from our companion paper: Wadekar et al. (2023a). For each mode template, we use a combination of post-Newtonian formulae and machine learning tools to model normalized aligned-spin templates corresponding to the and modes. We then filter the data separately with the three templates and store the complex SNR timeseries , see figure 1 of Wadekar et al. (2023a). As a result, our HM search matched-filtering cost is just that of a -only search (in contrast to being as in previous HM search methods, where the different modes are combined and templates sample over inclination and initial orbital phase Chandra et al. (2022); Schmidt et al. (2023); Harry et al. (2018)).
We divide our search space into 17 template banks, which are divided according to the shape of normalized waveforms corresponding to the binary parameters (the shape of normalized waveforms is roughly related to the binary merger frequency). The parameter space corresponding to each bank is shown in the bottom panel of Fig. 7 in the Appendix. Some of the banks are further divided into subbanks based on the range of chirp masses they cover. Note that the division into banks and subbanks is only to facilitate covering the parameter space with templates of the form that we use, and the final ranking statistic we assign to each trigger is weighed against those of all other triggers (over the entire parameter space) to estimate its global significance. Other than adding HM, we also improved our template banks by using a machine learning tool called random forest regressor (which helps us reduce the dimensionality of our template banks and enables us to search with 3040% fewer templates than before) Wadekar et al. (2023a).
@@ II.2 New statistics for combining SNR timeseries from different harmonics
We present the techniques for combining the timeseries in our companion paper (Wadekar et al. (2023b)) and here we provide a brief overview. Combining SNR timeseries of different modes incorrectly can lead to a large loss of sensitivity of the pipeline Capano et al. (2014). We use the Neyman–Pearson lemma to develop an optimal statistic to combine the SNR timeseries for different harmonics; our detection statistic is the ratio of the Bayesian evidence under the signal and the noise hypothesis. Ideally, one would marginalize over all the intrinsic and extrinsic binary parameters to calculate the evidence under the signal hypothesis. We develop two different approximate versions of the statistic to be used in different parts of our pipeline as detailed in the sub-sections below.
II.2.1 Single-detector marginalized statistic
We first developed an approximate statistic to collect individual detector triggers. Similar to other pipelines, we set a particular SNR threshold () to collect triggers from individual detectors. The primary reason behind this is to reduce the number of triggers passed to the next steps of the pipeline (coincidence analysis, veto and coherent score calculation) as these steps are computationally expensive (another reason is storage space constraints when saving a large number of triggers on disk). For collecting the triggers in a -only search, one can simply use . For the HM case, if we naively use , the number of collected triggers increases by a factor of over (using the same value of as the -only case Wadekar et al. (2023b)). To ameliorate this, we develop in Ref. Wadekar et al. (2023b) a new marginalized statistic () which exploits the fact that the values of and are bounded for physical signals (as opposed to Gaussian noise triggers where the values and can be arbitrarily large). The statistic includes marginalization over physically possible amplitude ratios of higher harmonics. To make the computation of faster, we maximize (instead of marginalizing) over the phases of HM and also over all the remaining intrinsic and extrinsic binary parameters (although this can be improved in future studies). We collect single-detector triggers using and we use a larger value of the collection threshold as compared to the -only case. The downside of increasing the value of is that low-SNR events can fall below our threshold bar. This is indeed the reason behind our pipeline missing some of the low-SNR events reported by -only pipelines in the previous literature, see Section III.3 below for more details. The particular choice of that we made in the current paper is currently not a bottleneck for our search and can be improved in the future.
@@ II.2.2 Coherent marginalized statistic
Once we collect single detector triggers, we check if the same template(s) fired in other detectors at a similar time. To rank the resulting coincident triggers (for calculating their false-alarm rate (FAR) and subsequently ), we develop a more accurate statistic which also includes coherence between different detectors for ranking the events. For example, we exploit the fact that astrophysical signals have correlated phase and time differences between the detectors, while the noise triggers have random distributions Nitz et al. (2017). In the previous sub-section, we maximized over extrinsic binary parameters, but marginalizing over them is the more optimal strategy. Our coherent ranking statistic thus includes marginalization over extrinsic parameters in a similar way to the IAS -only search methodology given in Appendix D of Olsen et al. (2022) (hereafter O22) except generalized to account for higher modes. In addition, we also include marginalization over the amplitude of HM (which effectively corresponds to marginalizing over mass ratios); the only parameters that we currently do not marginalize over are the intrinsic binary parameters corresponding to the 22 waveform, as we use only the best-fit 22 template in our calculations. Details regarding our coherent statistic will be given in upcoming papers Wadekar et al. (2023b); Roulet et al. (2023).
@@ II.3 Removing/downweighting noise transients
In addition to the changes involving higher modes, we also improve the sensitivity of our search for high masses compared to previous IAS pipelines by developing new tools. We also discuss in this sub-section methods to tailor the background reduction techniques in our -only pipeline for the case of HM.
II.3.1 Band eraser tool for data preprocessing
In our -only pipeline, in a data pre-processing step (before we matched-filter with templates), we identify regions of bad data segments which have excess power localized to particular frequency ranges and timescales (see section C of Venumadhav et al. (2019)). We then remove the data in the flagged regions (i.e., make “holes”) and then fill the holes using an “inpainting filter” (see figure 6 of Zackay et al. (2021)). Note that the holes remove the entire data from a particular time chunk. Sometimes, however, the noise transients are only localized to particular frequency bands (particularly the low-frequency end 100 Hz) and do not affect astrophysical signals which dominantly have power in a different frequency band.
Therefore, before making holes in entire time-chunks, we try to smoothly remove bad time-frequency regions (or bands) in the spectrogram. For this, we develop a new tool we call band eraser. We divide the spectrogram into multiple bands (each with dimensions ). We further split each band into multiple chunks (of ) and measure the number of chunks which have power above a certain threshold. If this number is above a certain threshold, we remove the entire band from our analysis. We calculate this threshold using Poisson statistics, making sure that, on average, in Gaussian noise, one band is removed per s (we also add an additional safety factor to the threshold to avoid removing gravitational wave signals).
We also noticed that some of the noise transients had power spread over long timescales ( s) and thus were sometimes missed by the localized excess-power tests (with timescales s) that we had used earlier to identify regions for making holes (see Table II of Venumadhav et al. (2019)). Empirically, we noticed that such disturbances predominantly pollute the low-frequency ( Hz) end, and therefore particularly reduced the sensitivity of our pipeline to high-mass mergers. We find that using the band eraser also helped in mitigating some such disturbances. We leave further discussion of the band eraser tool to Ref. Wadekar et al. (2023b).
II.3.2 Veto procedure
@@
In previous quadrupole-only searches, we reduced the noise background by accounting for the non-stationarity of the PSD Zackay et al. (2021). In the procedure, we match-filtered a reference template from the bank with the data and registered its SNR . We then estimated the variation in PSD as a function of time by averaging the power within rolling windows of length s. In this search, we calculate the correction for the PSD drift in the same way as before by only using waveforms from our template bank (and not including HM waveforms). In the limit in which the fluctuations of the PSD are described by a scalar over all frequencies, this procedure returns an unbiased estimate of the drift correction.
II.3.3 Ranking statistic

### v2

@@ Abstract
Nearly all of the previous gravitational wave (GW) searches in the LIGO–Virgo data included GW waveforms with only the dominant quadrupole harmonic, i.e., omitting higher-order harmonics which are predicted by general relativity. We improved the IAS pipeline by efficiently introducing higher harmonics in the GW templates using the techniques in Wadekar et al. (2023, 2024). Using the IAS-HM111The pipeline is publicly available at https://github.com/JayWadekar/gwIAS-HM pipeline on the public LIGO–Virgo data from the O3 run, we find 11 new candidate BBH mergers with (we use the detection threshold as the astrophysical probability, , being over 0.5, following the approach of other pipelines). We broadly recover the high-significance events from earlier catalogs, except a few which were vetoed. We also find that including higher harmonics in our search raises the significance of a few previously reported marginal events (e.g., GW190711_030756).
@@ I Introduction
Gravitational waves (GW) from roughly 100 binary black hole (BBH) mergers have been detected in the data from the first three observing runs of the LIGO–Virgo–Kagra collaboration (LVK), combining detections by the LVK Abbott et al. (2016a, 2019, 2021); The LIGO Scientific Collaboration et al. (2021); Abbott et al. (2021, 2022); Kumar and Dent (2024) and by other independent pipelines Venumadhav et al. (2019, 2020); Olsen et al. (2022); Nitz et al. (2019, 2020, 2021, 2021); Chia et al. (2024); Mehta et al. (2023); Koloniari et al. (2024) who access the data made publicly available through the Gravitational Wave Open Science Center (GWOSC) Vallisneri et al. (2015). Nearly all the previous template-bank based searches have approximated GW waveforms to have only the dominant quadrupole mode, , where the frequency of GWs is twice the orbital frequency.
@@
The mode appears at the lowest order in post-Newtonian (PN) expansion, but there are a number of higher harmonics originating from the subsequent orders of the PN expansion (e.g., the mode has thrice the orbital frequency) Thorne (1980); Pan et al. (2011). Approximate formulae for the relative strength of frequency-domain amplitudes of two of the higher-order modes (HM) are given by Wadekar et al. (2023)
@@
where is the mass ratio (), is the orbital frequency and is the binary inclination ( corresponds to the edge-on configuration). These higher harmonics have a number of important applications for characterizing properties of BBHs. If one uses the mode waveform alone in inferring the BBH properties, we have the following important degeneracies between different binary parameters: luminosity distance and ; and effective spin (); polarization and initial orbital phase (). All three degeneracies are broken by HM Mills and Fairhurst (2021); Varma et al. (2014). This follows from the amplitudes of HM in Eq. (1) having a different dependence on and as compared to that of the mode, and similarly the phases of HM have different dependence on . It is worth mentioning that the degeneracy between distance and inclination is currently one of the most important limiting factors in inferring the Hubble parameter from GW systems and HM can be crucial in breaking this degeneracy. Similarly, events which have HM also have the best measured mass ratios (e.g., Abbott et al. (2020a, b)). HM become important in the strong gravity regime close to merger and are thus useful to test the multipolar structure of gravity Kastha et al. (2018, 2019); Dhanpal et al. (2019); Islam et al. (2020); Capano et al. (2021); Islam et al. (2021a, b), and to study properties of the merger remnant such as the recoil kick Calderón Bustillo et al. (2018).
@@
Multiple waveform models including HM have been recently made available (e.g., London et al. (2018); Garc´ıa-Quirós et al. (2020); Mehta et al. (2017); Ramos-Buades et al. (2023)), and they are currently being used ubiquitously in parameter estimation studies. However, HM are still not included in current search pipelines (as of now, the only exception is Chandra et al. (2022), which however only searched for binaries with nearly edge-on inclinations). The reason behind this is that for making template banks with the mode, one only needs to sample over the masses and spins of black holes (). However, for constructing banks including HM, one needs to additionally sample over the inclination and initial phase (, ). These additional degrees of freedom can cause: the size of the template banks and the cost of search to increase significantly (by a factor of 100 for the case of HM Harry et al. (2018); Chandra et al. (2022); Schmidt et al. (2023)), loss of sensitivity in the search as there is increased number of background triggers Capano et al. (2014). To tackle the first issue, we found more efficient ways of including HM in our template banks (which we presented in detail in our first companion paper Wadekar et al. (2023)). Regarding the second issue, we devise new detection statistics which includes marginalization over the amplitudes and phases of higher modes in our second companion paper Wadekar et al. (2024). We present an overview of these new techniques in Section II.
@@
We see from Eq. (1) that, at a given orbital frequency, the relative HM amplitudes increase as a power-law with . Moreover, due to the shape of the current detector power spectral density (PSD), the signal-to-noise ratio (SNR) contribution of HM is additionally boosted at high masses and redshifts. In this regime, the lower-frequency mode can cut off below the detector’s sensitive band (the detector PSD sharply increases below Hz) and HM (being at higher frequencies) naturally have an increased contribution to the SNR. The background corresponding to the high-mass regime is however heavily dominated by short-duration noise transients (called instrumental “glitches”). This is because the high-mass waveforms have a very short duration in band, which makes it easy for transients like blip glitches to mimic real signals. We therefore introduce additional methods in the IAS pipeline to downweight the background at high masses. Some of these methods are detailed in our second companion paper Wadekar et al. (2024) (hereafter W24) and we present an overview in Section II.3.
@@ II Pipeline improvements
Here, we present an outline of the additions that we made to the previous -only version of the IAS pipeline Olsen et al. (2022); Venumadhav et al. (2019).
@@ II.1 Adding HM in template banks
In this sub-section, we summarize the methodology for constructing template banks for HM from our companion paper: Wadekar et al. (2023). For each mode template, we use a combination of post-Newtonian formulae and machine learning tools to model normalized aligned-spin templates corresponding to the and modes. We then filter the data separately with the three templates and store the complex SNR timeseries , see figure 1 of Wadekar et al. (2023). As a result, our HM search matched-filtering cost is just that of a -only search (in contrast to being as in previous HM search methods, where the different modes are combined and templates sample over inclination and initial orbital phase Chandra et al. (2022); Schmidt et al. (2023); Harry et al. (2018)).
We divide our search space into 17 template banks, which are divided according to the shape of normalized waveforms corresponding to the binary parameters (the shape of normalized waveforms is roughly related to the binary merger frequency). The parameter space corresponding to each bank is shown in the bottom panel of Fig. 7 in the Appendix. Some of the banks are further divided into subbanks based on the range of chirp masses they cover. Note that the division into banks and subbanks is only to facilitate covering the parameter space with templates of the form that we use, and the final ranking statistic we assign to each trigger is weighed against those of all other triggers (over the entire parameter space) to estimate its global significance. Other than adding HM, we also improved our template banks by using a machine learning tool called random forest regressor (which helps us reduce the dimensionality of our template banks and enables us to search with 3040% fewer templates than before), see Fig. 5 of Wadekar et al. (2023).
@@ II.2 New statistics for combining SNR timeseries from different harmonics
We present the techniques for combining the timeseries in our companion paper (W24) and here we provide a brief overview. Combining SNR timeseries of different modes incorrectly can lead to a large loss of sensitivity of the pipeline Capano et al. (2014). We use the Neyman–Pearson lemma to develop an optimal statistic to combine the SNR timeseries for different harmonics; our detection statistic is the ratio of the Bayesian evidence under the signal and the noise hypothesis Neyman and Pearson (1933). Ideally, one would marginalize over all the intrinsic and extrinsic binary parameters to calculate the evidence under the signal hypothesis. We develop two different approximate versions of the statistic to be used in different parts of our pipeline as detailed in the sub-sections below.
II.2.1 Single-detector collection statistic
We first developed an approximate statistic to collect individual detector triggers. Similar to other pipelines, we set a particular SNR threshold () to collect triggers from individual detectors. The primary reason behind this is to reduce the number of triggers passed to the next steps of the pipeline (coincidence analysis, veto and coherent score calculation) as these steps are computationally expensive (another reason is storage space constraints when saving a large number of triggers on disk). For collecting the triggers in a -only search, one can simply use . For the HM case, if we naively use , the number of collected triggers increases by a factor of over (using the same value of as the -only case (W24)). To ameliorate this, we develop a new marginalized statistic to collect triggers ( in W24). We exploit the fact that the values of and are bounded for physical signals (as opposed to Gaussian noise triggers where the values and can be arbitrarily large), see Fig. 2 of (W24). We therefore marginalize over the space of physically possible amplitude ratios of higher harmonics using a pre-computed prior samples for (which are derived from our astrophysical prior). Note that the prior on can be very different for templates in different regions of parameter space. For example, the median of our prior for is 0.04 for Bank 0 (a low-mass bank), while it is 0.3 for Bank 14 (a high-mass bank). If we did not include such a prior, we would indeed expect to encounter a large loss of sensitivity on including HM. To make the computation of faster, we maximize (instead of marginalizing) over the phases of HM and also over all the remaining intrinsic and extrinsic binary parameters.
@@ II.2.2 Coherent marginalized statistic
Once we collect single detector triggers, we check if the same template(s) fired in other detectors at a similar time. To rank the resulting coincident triggers (for calculating their false-alarm rate (FAR) and subsequently ), we develop a more accurate statistic which also includes coherence between different detectors for ranking the events. For example, we exploit the fact that astrophysical signals have correlated phase and time differences between the detectors, while the noise triggers have random distributions Nitz et al. (2017). In the previous sub-section, we maximized over extrinsic binary parameters, but marginalizing over them is the more optimal strategy. Our coherent ranking statistic thus includes marginalization over extrinsic parameters in a similar way to the IAS -only search methodology given in Appendix D of Olsen et al. (2022) (hereafter O22), except generalized to account for higher modes (see Roulet et al. (2024); Leslie et al. (2021) for details). In addition, we also include marginalization over the amplitude of HM (which effectively corresponds to marginalizing over mass ratios); the only parameters that we currently do not marginalize over are the intrinsic binary parameters corresponding to the 22 waveform, as we use only the best-fit 22 template in our calculations. Details regarding our coherent statistic are provided in our companion papers Wadekar et al. (2024); Roulet et al. (2024).
@@ II.3 Removing/downweighting noise transients
In addition to the changes involving HMs, we also improve the sensitivity of our search for high masses compared to previous IAS pipelines by developing new noise reduction tools.
II.3.1 Trigger Inference using Extended strain representation (TIER)
We developed a machine learning (ML) framework called TIER (Trigger Inference using Extended strain Representation) for improving the sensitivity of any generic pipeline Wadekar et al. (a). Typically, search pipelines only use a small region of strain data in the vicinity of a candidate signal to construct the detection statistic. However, extended strain data ( s) in the candidate’s vicinity can also carry valuable complementary information. We included this extended information into our detection statistic using ML classifiers (see Fig. 1 of Wadekar et al. (a)). Using TIER, we found up to 20% improvement in sensitive volume of our pipeline, with improvements being concentrated in regions of high masses and unequal mass ratios. We refer the reader to Wadekar et al. (a) for further details regarding TIER.
II.3.2 Band eraser tool for data preprocessing
In our -only pipeline, in a data pre-processing step (before we matched-filter with templates), we first identify regions of noisy data segments (see section C of Venumadhav et al. (2019)), and remove the data in the flagged regions (i.e., make “holes”). We then fill the holes using an “inpainting filter” (see Fig. 6 of Zackay et al. (2021)). Note that the holes remove the entire data from a particular time chunk. Sometimes, however, the noise transients are only localized to particular frequency bands (particularly the low-frequency end 100 Hz) and do not affect astrophysical signals which dominantly have power in a different frequency band.
Therefore, before making holes in entire time-chunks, we try to smoothly remove bad time-frequency regions (or bands) in the spectrogram by developing a new tool called band eraser. First, we divide the spectrogram into multiple bands (each with dimensions ). We further split each band into multiple chunks (of ) and measure the number of chunks which have power above a certain threshold. If this number of chunks in a given band is above a certain threshold, we remove the entire band from our analysis. We leave further discussion of the band eraser tool to W24.
II.3.3 Improving signal consistency vetos using ML
@@
One of the downsides of our veto tests is that if a loud signal displays effects which are not currently present in our templates (e.g. precession or eccentricity), the chances of a veto test mis-identifying it as a glitch are much higher (e.g. GW190521 was a high-mass merger which got vetoed in a previous IAS search O22 as it shows a mild evidence for precession). To remedy this, we run our pipeline on precessing GW signals injected in the data. We then use a ML classifier to distinguish between properties of GW signals which got vetoed, as opposed to the properties of the noise transients which got vetoed. The ML classifier therefore identifies the precise combination of signal consistency tests which are more likely to fire when a GW signal is precessing. The ML classifier output probability is then used to set a criteria for overriding the vetos in a select few cases. This helps us recover loud precessing signals and increases the volume-time sensitivity of our pipeline by 10% across nearly all of the parameter space. We leave further discussion of the ML-based veto overriding procedure to a future paper Wadekar et al. (a).
@@
In previous quadrupole-only searches, we reduced the noise background by accounting for the non-stationarity of the PSD Zackay et al. (2021). In this procedure, we matched-filter a reference template from the bank with the data and register its SNR . We then estimate the variation in PSD as a function of time by averaging the power within rolling windows of length s. In the limit in which the fluctuations of the PSD over all frequencies are described by a single scalar, this procedure returns an unbiased estimate of the drift of the PSD. In our HM search, we calculate the correction for the PSD drift in the same way as before by only using a reference waveform from the template bank for calculating . We currently do not include (3,3) or (4,4) waveforms in estimating PSD drift and leave this direction to future work.
II.3.4 Ranking statistic